# Quantum Mycielskians: symmetries, twin vertices and distinguishing labelings

### v1

@@ Acknowledgments
$A_{G_{1}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{2}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{3}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&1&1&0&1\\ 1&1&0&0&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{4}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{5}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{6}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&0\\ 0&0&1&1&0&1\\ 1&1&0&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{7}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&1\\ 0&0&1&1&0&1\\ 1&1&0&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{8}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&0&1&1&0&0\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{9}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&1&1&1&0&0\\ 1&0&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{10}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&0&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{11}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&1&1&0&0\\ 1&0&1&1&0&0\end{bmatrix},\quad\det=-1,$ $A_{G_{12}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&1&1&1&0\end{bmatrix},\quad\det=-1,$
$A_{G_{13}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 0&1&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=-1,$ $A_{G_{14}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&0&1&1&0\\ 0&1&1&0&0&1\\ 0&1&1&0&0&1\\ 1&0&0&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{15}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&0&1&1&0\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&0&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{16}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&1\\ 0&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{17}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&0\\ 0&1&1&1&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{18}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{19}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&1&1&1&0\\ 0&1&0&1&1&0\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=3,$ $A_{G_{20}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$
$A_{G_{21}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{22}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{23}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{24}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{25}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{26}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&1&0\\ 1&1&0&1&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{27}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{28}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{29}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&0\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{30}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{31}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{32}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{33}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&0&1\\ 1&0&0&0&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=3,$ $A_{G_{34}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&1&1\\ 1&0&0&1&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=4,$
$A_{G_{35}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&0&1&1&0&1\\ 1&0&1&1&1&0\end{bmatrix},\quad\det=4,$ $A_{G_{36}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&0&1\\ 1&0&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=7,$
$A_{G_{37}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&0\\ 1&0&0&1&0&0\\ 1&1&1&0&0&0\end{bmatrix},\quad\det=-5,$ $A_{G_{38}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&1\\ 1&0&0&1&0&0\\ 1&1&1&1&0&0\end{bmatrix},\quad\det=0,$
$A_{G_{39}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{40}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=4,$
$A_{G_{41}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{42}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{43}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&1\\ 1&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{44}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$\ A_{G_{45}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&0&1&1\\ 1&0&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{46}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&0&1&1\\ 1&0&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=7,$
$A_{G_{47}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&1&1&1\\ 1&0&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{48}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&1&1&1\\ 1&0&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$
$A_{G_{49}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix},\quad\det=-1,$ $A_{G_{50}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&0&1\\ 1&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{51}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$ $A_{G_{52}^{6}}$ $=\begin{bmatrix}0&0&1&1&1&1\\ 0&0&1&1&1&1\\ 1&1&0&0&1&1\\ 1&1&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$
[…]

### v2

@@ Acknowledgments
$A_{\scriptscriptstyle G_{1}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{2}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{3}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&1&1&0&1\\ 1&1&0&0&1&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{4}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{5}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{6}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&0\\ 0&0&1&1&0&1\\ 1&1&0&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{7}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&1\\ 0&0&1&1&0&1\\ 1&1&0&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{8}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&0&1&1&0&0\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{9}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&1&1&1&0&0\\ 1&0&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{10}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&0&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{11}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&1&1&0&0\\ 1&0&1&1&0&0\end{bmatrix},\quad\det=-1,$ $A_{\scriptscriptstyle G_{12}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&1&1&1&0\end{bmatrix},\quad\det=-1,$
$A_{\scriptscriptstyle G_{13}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 0&1&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=-1,$ $A_{\scriptscriptstyle G_{14}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&0&1&1&0\\ 0&1&1&0&0&1\\ 0&1&1&0&0&1\\ 1&0&0&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{15}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&0&1&1&0\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&0&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{16}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&1\\ 0&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{17}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&0\\ 0&1&1&1&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{18}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{19}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&1&1&1&0\\ 0&1&0&1&1&0\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=3,$ $A_{\scriptscriptstyle G_{20}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$
$A_{\scriptscriptstyle G_{21}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{22}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{23}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{24}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{25}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{26}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&1&0\\ 1&1&0&1&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{27}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{28}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{29}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&0\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{30}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{31}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{32}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{33}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&0&1\\ 1&0&0&0&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=3,$ $A_{\scriptscriptstyle G_{34}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&1&1\\ 1&0&0&1&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=4,$
$A_{\scriptscriptstyle G_{35}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&0&1&1&0&1\\ 1&0&1&1&1&0\end{bmatrix},\quad\det=4,$ $A_{\scriptscriptstyle G_{36}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&0&1\\ 1&0&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=7,$
$A_{\scriptscriptstyle G_{37}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&0\\ 1&0&0&1&0&0\\ 1&1&1&0&0&0\end{bmatrix},\quad\det=-5,$ $A_{\scriptscriptstyle G_{38}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&1\\ 1&0&0&1&0&0\\ 1&1&1&1&0&0\end{bmatrix},\quad\det=0,$
$A_{\scriptscriptstyle G_{39}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{40}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=4,$
$A_{\scriptscriptstyle G_{41}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{42}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{43}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&1\\ 1&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{44}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$\ A_{\scriptscriptstyle G_{45}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&0&1&1\\ 1&0&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{46}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&0&1&1\\ 1&0&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=7,$
$A_{\scriptscriptstyle G_{47}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&1&1&1\\ 1&0&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{\scriptscriptstyle G_{48}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&1&1&1\\ 1&0&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$
$A_{\scriptscriptstyle G_{49}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix},\quad\det=-1,$ $A_{\scriptscriptstyle G_{50}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&0&1\\ 1&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{\scriptscriptstyle G_{51}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$ $A_{\scriptscriptstyle G_{52}^{6}}$ $=\begin{bmatrix}0&0&1&1&1&1\\ 0&0&1&1&1&1\\ 1&1&0&0&1&1\\ 1&1&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$
[…]